### v1

# Enhancing Repository-Level Code Generation with Integrated Contextual Information

Abstract.
Large language models (LLMs) have demonstrated remarkable capabilities in code generation tasks. However, repository-level code generation presents unique challenges, particularly due to the need to utilize information spread across multiple files within a repository. Existing retrieval-based approaches sometimes fall short as they are limited in obtaining a broader and deeper repository context. In this paper, we present CatCoder, a novel code generation framework designed for statically typed programming languages. CatCoder enhances repository-level code generation by integrating relevant code and type context. Specifically, it leverages static analyzers to extract type dependencies and merges this information with retrieved code to create comprehensive prompts for LLMs. To evaluate the effectiveness of CatCoder, we adapt and construct benchmarks that include 199 Java tasks and 90 Rust tasks. The results show that CatCoder outperforms the RepoCoder baseline by up to 17.35%, in terms of pass@$k$ score. Furthermore, the generalizability of CatCoder is assessed using various LLMs, including both code-specialized models and general-purpose models. Our findings indicate consistent performance improvements across all models, which underlines the practicality of CatCoder.
Large Language Model, Code Generation, Repository Context
copyright: none
Figure 1. A scenario of developing a Java method triu in the Apache Commons Math (com, 2024) library. (1) The left part illustrates relevant code, whose patterns might be helpful when developers write current method code. (2) The right part illustrates the type context (i.e., the fields and method signatures of a class) related to the current method, which can be derived iteratively through static analysis, starting with the method signature.
1. Introduction
Large language models (LLMs) (Wang et al., 2021; Nijkamp et al., 2022; Roziere et al., 2023; Lozhkov et al., 2024) have been widely applied in various software engineering-related fields, and they have shown remarkable performance in code generation tasks (Chen et al., 2021; Austin et al., 2021). Code generation tasks can be divided into two types: standalone code generation and repository-level code generation. For the standalone scenario, the function can be invoked independently, and the generation of LLMs is conditioned only on the in-file context (i.e., contents within the same file). However, for more practical repository-level code generation, LLMs struggle to generate correct code if only in-file context is provided, due to that relevant information usually lies in many different files in a repository, and the lack of cross-file context is likely to lead to hallucinations. Therefore, it is essential to fully utilize the repository information and provide cross-file contexts for the LLMs to generate the correct repository-level code.
To utilize repository information, existing studies use retrieval-based approaches to find similar code for reference. Zan et al. (Zan et al., 2022) train a dual-encoder model to retrieve API information from library documentation. Lu et al. (Lu et al., 2022) build an external code database to retrieve similar code. Shrivastava et al.  (Shrivastava et al., 2023) propose RLPG, in which they train a classifier to predict prompt proposals taken from repository files. Zhang et al. (Zhang et al., 2023) propose RepoCoder, which iteratively retrieves similar code in a repository. However, these approaches still have limitations. First, for approaches that use external sources or train models, they are not generalizable to new unseen repositories. Second, retrieved code cannot always cover the necessary logic and API invocations that correct code requires, especially when a code repository does not contain much duplicated code (Zhang et al., 2023).
To address the latter limitation, we find that apart from code retrieval, in statically typed programming languages [^1] Statically typed programming languages are those where the data types of variables have to be declared during compile-time, such as C, C++, Java and Rust., the types related to a function can also provide valuable references to LLMs. In other words, they can provide a set of fields and methods for the LLM to select and use in the generated code. In this paper, we refer to this information as type context. The type context serves as a complement to the code retrieval and helps LLMs better understand the repository structure at a deep level. To extract such information, we utilize existing static analyzers.
Based on type context, we propose a new repository-level code generation framework named CatCoder, which exploits the combination of relevant code and type context for statically typed languages. It works on top of a frozen LLM and does not require external databases or model training. First, it retrieves relevant code from the repository. Then, CatCoder invokes a static analyzer to obtain dependent types for the function and creates textual representations of the type context. Finally, CatCoder merges two kinds of information into the prompt and generates the function code with an LLM.
To evaluate the effectiveness of CatCoder, we adapt and construct repository-level code generation benchmarks that include 199 Java tasks and 90 Rust tasks in total. We compare CatCoder with vanilla LLM, the In-File context, and RepoCoder (Zhang et al., 2023), a state-of-the-art framework for repository-level code generation, in terms of pass@$k$ (Chen et al., 2021) scores. To validate the generalizability of CatCoder, we evaluate CatCoder with various LLMs, including code-specialized models and general purpose models. The results on the benchmarks indicate that CatCoder substantially outperforms the vanilla LLM and the In-File context. In addition, it outperforms RepoCoder by up to 17.35%. The results of the generalizability study demonstrate that CatCoder improves the performance of all selected LLMs.
In summary, we make the following contributions:
  1. (1)
    We propose a novel framework for statically-typed programming languages, CatCoder, to generate repository-level code by highlighting relevant code and type context in code repositories.
  2. (2)
    To the best of our knowledge, we are the first to construct a repository-level code generation benchmark for Rust, to evaluate the effectiveness of CatCoder on a minor programming language.
  3. (3)
    We evaluate CatCoder using Java and Rust benchmarks, and results indicate that CatCoder outperforms the baselines.
Paper Organization: The remainder of the paper is structured as follows. Section 2 introduces the motivation of CatCoder using an example. Section 3 presents the details of CatCoder. Section 4 describes the experimental setups, including baselines, evaluation metric, benchmarks, and implementation details. Section 5 analyzes the experimental results and provides answers to the research questions. Section 6 discusses the failure cases of CatCoder and threats to validity. Section 7 summarizes the related work. Section 8 concludes the paper with possible future work.
Figure 2. Overview of our approach
2. Motivating Example
In this section, we present a motivating example to illustrate the main idea of CatCoder. Figure 1 shows a scenario for developing a Java method named triu|, which takes a RealMatrix object and an int as input and returns a RealMatrix object.
From the perspective of human developers, to write repository-level code, we do not only refer to the method signature and the docstrings but also search for relevant code as references. In addition, for statically typed programming languages like Java, it is common to dive into related types (e.g., argument and return types) to seek available APIs. The two kinds of information can be concluded as follows:
  1. Info.1
    Relevant code with a similar name, parameters, return type, or docstrings. As similar code is more likely to perform similar or symmetric operations, we can learn from the patterns and write new code.
  2. Info.2
    Type context, i.e., the field and method information of related classes, which provides hints on the possible ways to assign values or invoke APIs.
[…]
After obtaining such information, developers can write the code by selecting useful parts of the relevant code and the type context and combining them with some logic. Given the powerful in-context learning ability of LLMs, it is intuitive that LLMs can also benefit from such comprehensive information if we utilize it to augment the prompts.
To improve the performance of repository-level code generation with LLMs, existing work (Lu et al., 2022; Zhang et al., 2023) uses retrieval-augmented generation to retrieve relevant code as presented in Info.1. However, it is limited as it ignores Info.2 (i.e., type context). For code repositories that do not contain many similar method pairs, the retrieved code would be less helpful. In this case, without type context, LLMs are more likely to hallucinate and generate non-existent field accesses or API calls.
Motivated by the aforementioned example, we illustrate the idea of our work CatCoder, which highlights the combination of Info.1 and Info.2. CatCoder is carried out in the following steps:
  1. Step a.
    Relevant Code Retrieval. We retrieve similar code in a repository based on the query (i.e., the docstrings and signature of the function to be completed).
  2. Step b.
    Type Context Extraction. We utilize static analyzers to determine the dependent types and then create type context by extracting the fields and method signatures.
  3. Step c.
    Code Generation via LLM. We merge the results in Step a and Step b into a comprehensive prompt, and invoke an LLM to obtain the generated code.
3. Proposed Approach
The framework of CatCoder is presented in Figure 2. CatCoder takes a source code repository and an empty function (along with its docstrings, if it exists). The output of CatCoder is the completed code of the function.
In addition to Figure 2, Algorithm 1 presents the pseudocode of the workflow of CatCoder. CatCoder contains two core components: (1) Relevant Code Retrieval and (2) Type Context Extraction. By utilizing these components, CatCoder is able to gather two kinds of referential information for the LLM. After post-processing the raw output of the LLM, CatCoder returns the generated code.
Input: Code repository $R$, Empty function $f$
Output: The completed code of function $f$
1
2
docs = code_retrieval($R,f$)
  /* Section 3.1 */
contexts = type_context($R,f$)
  /* Section 3.2 */
prompt = prepare($f$, docs, contexts)
  /* Section 3.3 */
3 code = LLM(prompt)
4 return postprocess(code)
Algorithm 1 CatCoder
3.1. Relevant Code Retrieval
3.1.1. Code Splitting
We traverse all source code files in the repository and split them into smaller chunks. Starting with complete files, we set a maximum chunk size and recursively split larger fragments into smaller ones that fit in the chunk.
As the structure of the source code is often complex, we cannot split it at an arbitrary position. Otherwise, the resulting contents would be less meaningful and difficult to understand. Therefore, to better preserve the code structures, we split the code only before the following positions (sorted by priority in descending order):
  1. (1)
    Type definitions
  2. (2)
    Function definitions
  3. (3)
    Control flow statements
  4. (4)
    New lines
After splitting, we obtain a corpus of code chunks [^2] We call them “documents” later in this subsection. to perform retrieval.
3.1.2. Hybrid Retrieving
[…]
❶ Sparse Retrieval
We choose bag-of-words model for sparse retrieval, and use the BM25 function (Robertson et al., 1995; Trotman et al., 2012) to rank the set of documents, based on the query terms. Given a document $D$ and a query $Q$ with terms $t_{1},t_{2},\cdots,t_{n}$, the BM25 score can be defined as follows:
[…]
where $\mathrm{IDF}(t,D)$ is the inverse document frequency weight, $\mathrm{TF}(t,D)$ is the term frequency (i.e., the number of appearing times in the document), $|D|$ is the length of the document, and $k_{1}$ and $b$ are hyperparameters. In default, the values of $k_{1}$ and $b$ are set to 1.5 and 0.75. $\mathrm{IDF}(t,D)$ can be calculated by:
$\mathrm{IDF}(t,D)=\log\dfrac{N_{D}-\mathrm{DF}(t,D)+0.5}{\mathrm{DF}(t,D)+0.5},$
where $N_{D}$ is the number of documents and $\mathrm{DF}(t,D)$ is the document frequency (i.e., the number of documents that contain the term).
In practice, after splitting the code files from a repository, we apply the aforementioned BM25 function for a query and obtain a list of scores. After that, we select documents with top-$k$ highest scores.
❷ Dense Retrieval
We employ an embedding model for dense retrieval. The model maps natural language sentences and code snippets into dense vector spaces. To select the most relevant documents, our intuitive idea is to find the nearest neighbors around the embedding vector of the query text. Thus, we use Squared Euclidean Distance to measure the proximity between embedding vectors in high-dimensional spaces, and this metric can be calculated by:
$d^{2}(\mathbf{p},\mathbf{q})=\left\|\mathbf{p}-\mathbf{q}\right\|^{2}=\sum_{i=1}^{n}(p_{i}-q_{i})^{2},$
where $n$ is the number of dimensions in the embedding vectors. A lower value of $d^{2}(\mathbf{p},\mathbf{q})$ indicates that $\mathbf{p}$ and $\mathbf{q}$ are more similar.
In practice, we utilize the embedding model and generate vector representations of all documents and query text. Then we calculate $d^{2}(\mathbf{p},\mathbf{q})$ between the query embedding and all document embeddings and select the top-$k$ nearest documents.
3.1.3. Result Ranking
In the third step, we ensemble the two retrievers and re-rank their outputs using the Reciprocal Rank Fusion (RRF) (Cormack et al., 2009) algorithm. We assign weights to the retrievers and calculate the weighted RRF score [^4] We assume that a retriever does not output duplicate documents. However, duplicate documents across retrievers are acceptable. for each retrieved document $D$:
$\mathrm{RRF}(D)=\sum_{i=1}^{n}\mathbf{1}(D)\cdot\dfrac{w_{i}}{r(D)+60},$
where $w_{i}$ is the weight assigned to a certain retriever, $\mathbf{1}(D)$ is an indicator function [^5] The value is 1 if $D$ is one of the outputs of this retriever, otherwise the value is 0., and $r(D)$ is the original rank (if exists) of $D$ in the outputs of this retriever.
After applying the RRF algorithm, we deduplicate the documents and sort them in descending order based on the RRF scores. Now we obtain the results of the Relevant Code Retrieval component.
3.2. Type Context Extraction
Here, we present the type context and the way to extract type contexts from code repositories using static analyzers [^6] The word “type” is an abstract concept and might be different in various languages. For example, it refers to class, interface and enum in Java; whereas in Rust, it refers to struct and enum..
3.2.1. Type Context
Figure 3. The dependency graph of the Java class CMAESOptimizer
Figure 3 shows an undirected dependency graph $G=(V,E)$, in which each vertex corresponds to a type and each edge represents the dependency relationship. In particular, if type B occurs in the implementation of type A, there should be an edge between A and B, and vice versa. Thus, the dependency graph contains types that we are likely to use and refer to when writing code of method body.
In Figure 3, we build the dependency graph starting from class CMAESOptimizer. As the graph spans, the vertex set and the edge set will eventually converge, and we can obtain the types that are relevant to CMAESOptimizer. To pass such type information to the LLM, we create type context from the graph by extracting the fields and function signatures within the types and concatenating them into a string.
However, in large code repositories, the number of vertices could grow exponentially, making the resulting type context too long to fit in an LLM’s context window. In addition, higher-order neighbor vertices in the graph tend to have less relevance to each other. For the example in Figure 3, programmers are more likely to use the API of RealMatrix rather than RealVector, which is a higher-order neighbor of CMAESOptimizer compared to RealMatrix. Therefore, in practice, we limit the scope of type context creation to one-order neighbors in the graph. Assuming that LLMs have knowledge of the standard libraries of many languages, we also prune the types in these libraries.
3.2.2. Building Type Context via Static Analyzer
To get detailed information (i.e., fields and methods) of a type from its name, some straightforward heuristics might be used, e.g., locating imports or looking for similar file names. However, these heuristics are not sound and do not work with all statically typed programming languages. Thus, we need to perform an analysis to accurately navigate to a type and find its definitions and implementations. Many language servers serve as static analyzers and provide this type-based analysis, such as clangd (cla, 2024) for C++, Eclipse JDT (ecl, 2024) for Java, and rust-analyzer (rus, 2024b) for Rust.
Specifically, to build the type context for a function code generation task, we first determine the set of relevant types (as Section 3.2.1 describes). We start with the signature and form a set of types, and the initial element is the type to which the current function belongs. Then we implement a parser that works collaboratively with the static analyzer. The static analyzer provides detailed type information to the parser. The parser then searches for a new occurring type, adds it to the set, and queries the analyzer with its location (i.e., byte offset or line and column offsets in a file). This interaction ends when all one-order neighbors of the initial types are found. Finally, we concatenate the fields and method signatures within the set of types and obtain the type context.
3.3. Code Generation via LLM
Figure 4. Our prompt design. The actual prompt ends right after the signature.
Finally, we describe how to utilize the previous results and invoke the LLM to generate code. Figure 4 illustrates our prompt design, in which there are two fields: CONTEXT and CODE. To construct a prompt, we fill the CONTEXT field with type context and retrieved code sequentially, and place the docstrings and the signature of the method at the beginning of the CODE field. The prompt serves as input to a frozen LLM [^7] We discuss the details of selected models and model access in Section 4.4., and we expect the model to complete the CODE field by generating the method body (i.e., the green part in Figure 4). In addition, to post-process the raw output of the LLM, we employ several procedures [^8] The post-processing procedures may vary with different LLMs and programming languages., including truncating texts, prepending function signatures, and removing Markdown syntaxes.
4. Experimental Setup
In this section, we first formulate three research questions (RQs). Based on the RQs, we describe the details of the baselines, evaluation metric, benchmarks, and implementation of our approach. We aim to answer the following research questions:
  1. RQ.1
    How effective is our approach at generating function code in a repository, compared with the baselines?
  2. RQ.2
    How do the two components in our approach (i.e., relevant code retrieval and type context extraction) contribute to the overall effectiveness respectively?
  3. RQ.3
    Is our approach generalizable to various large language models? And how do different models perform on the benchmark?
In brief, we evaluate the effectiveness of our approach in RQ1, and then conduct an ablation study in RQ2. Finally, in RQ3 we study the generalizability of our approach and the performance across several LLMs. For implementation and evaluation, we select two statically typed programming languages, Java and Rust, since Java is one of the most widely used programming languages (tio, 2024), and Rust is a new system-level language that attracts the attention of many developers (rus, 2024c).
4.1. Baselines
As described in Section 3.3, Figure 4 shows the prompt template used in our evaluation. As the CONTEXT field is the only variable, we compare our approach with methods that fill different contexts in the prompt. For the repository-level code generation task, there are three types of context: (1) No context, (2) Local context, and (3) Cross-file context. Therefore, we design the following baselines for comparison in RQ1:
  1. (1)
    Vanilla: In this method, we prompt the LLMs without any repository context. In this case, all natural language descriptions are given, but the CONTEXT field in Figure 4 is left empty.
  2. (2)
    In-File: In this method, we consider code generation as a left-to-right file completion task. We extract code before the function to be generated, and put it in the CONTEXT field.
  3. (3)
    RepoCoder: RepoCoder (Zhang et al., 2023) is an approach that augments the prompt by iteratively retrieving referential code fragments, and it is originally proposed for Python code repositories. We adopt their idea and implement this baseline for Java and Rust based on our retrieval pipeline. Specifically, the initial iteration uses natural language descriptions (i.e., docstrings) as the query for code retrieval, and the second iteration uses previously generated code for retrieval.
4.2. Evaluation Metric
For function code generation task, we argue metrics based on text similarity are not accurate, and it is essential to use metrics based on execution results. Therefore, to evaluate the functional correctness of the generated code, we follow Chen et al. (Chen et al., 2021) and use the unbiased pass@$k$ as a metric. This metric is defined as
[…]
where $n$ is the number of code samples generated per task, $c$ is the number of functional correct code samples, and $k$ is the parameter in “pass@$k$” ($n\geq k$). Here, a code sample is functionally correct if and only if it passes the tests in the benchmark or dataset. We calculate pass@$k$ values for each task, and the final pass@$k$ for a benchmark is the mean value of all pass@$k$ values.
4.3. Benchmarks
4.3.1. Java Benchmark
There are several publicly available repository-level benchmarks for Java (Ding et al., 2024; Yu et al., 2024; Agrawal et al., 2024; Zeng et al., 2024). However, it is challenging to use many of them to evaluate CatCoder because they either lack the raw repository code or are not designed for function code generation tasks. Considering the challenges, the CoderUJB (Zeng et al., 2024) benchmark is a suitable base benchmark that contains 238 code generation tasks, and we successfully adapt it for our evaluation. In the adaptation process, we filter out a few functions that satisfy the following patterns:
  1. (1)
    The length of function code exceeds max output limit (i.e., 512 tokens in our settings).
  2. (2)
    The function is standalone, which means that it can be moved to a new class [^9] We assume that all Java standard libraries are imported in advance when applying this rule. and still compiles. The removal of standalone functions helps better evaluate the performance of repository-level code generation.
  3. (3)
    The function does not have any modifiers and annotations. This helps us perform post-processing more easily.
After filtering, we finally obtain 199 coding tasks. As the methods in the benchmark are mined by the CoderUJB authors from the Defects4J (Just et al., 2014) dataset, it is convenient to test the correctness of the generated code utilizing the existing unit test methods. Specifically, we “patch” the repository with generated code and run the defects4j test -t ... command.
4.3.2. Rust Benchmark
To our knowledge, the only code generation benchmark for the Rust programming language is MultiPL-E (Cassano et al., 2023), which translates the original HumanEval (Chen et al., 2021) and MBPP (Austin et al., 2021) benchmarks. However, this benchmark is designed for standalone function code generation tasks and thus is unsuitable for our evaluation. In this case, we construct a new benchmark named RustEval.
Table 1. The crates used in our Rust benchmark
Crate # Downloads Crate # Downloads
bit-vec 54M sdl2 1.7M
curv-kzen 161K splay_tree 72K
dlv-list 14M sshkeys 189K
gptman 514K unix_path 86K
mathru 202K vec_map 81M
multiset 99K without-alloc 259K
rocket_http 4.2M
As shown in Table 1, we collect 13 repositories of various categories from crates.io (cra, 2024), the official Rust package registry. In these repositories, we obtain 90 functions with docstrings that satisfy the official guidelines (rus, 2024a) (i.e., have detailed descriptions and at least one documentation test). To test the correctness of the generated code, we utilize the documentation tests in the docstrings by running the cargo test --doc ... command.
4.4. Implementation
4.4.1. Approach
We implement CatCoder for Java and Rust, and the source code is written in Python and Rust. Specifically in Relevant Code Retrieval, we use MPNet V2 (mpn, 2024) as the embedding model. It is an open-source model fine-tuned from the base MPNet (Song et al., 2020) model using massive training data, including the popular CodeSearchNet (Husain et al., 2019). In Type Context Extraction, we use Eclipse JDT.LS (ecl, 2024) and rust-analyzer (rus, 2024b) as the static analyzers. For Java, our approach interacts with Eclipse JDT.LS through the Language Server Protocol (lsp, 2024). For Rust, we use rust-analyzer as a third-party library and implement a binding between Rust and Python.
4.4.2. Model Selection
Table 2. Models used for evaluation in RQ3. The model acronyms correspond to the $x$-axis labels in Figure 6.
Model ID Acronym Open-Source?
Code Specialized
(Roziere et al., 2023) CodeLlama-7b-Instruct CL-7B
(Roziere et al., 2023) CodeLlama-13b-Instruct CL-13B
(Roziere et al., 2023) CodeLlama-34b-Instruct CL-34B
(Guo et al., 2024) deepseek-coder-6.7b-instruct DS-6.7B
(cod, 2024) codegemma-7b-it CG-7B
(Bai et al., 2023) CodeQwen1.5-7B-Chat CQ-7B
(Lozhkov et al., 2024) starcoder2-15b SC-15B
General Purpose
(lla, 2024) Meta-Llama-3-8B-Instruct ML-8B
(gpt, 2024) gpt-3.5-turbo-0125 GPT-3.5
For RQ1 and RQ2, we perform the evaluation using CodeLlama-13B-Instruct (Roziere et al., 2023) as the default LLM, due to its state-of-the-art performance among models of medium size. For RQ3, to evaluate the generalizability of CatCoder, we select a variety of state-of-the-art LLMs with different sizes, including code-specialized models and general-purpose models. The details of the models are listed in Table 2, in which the model IDs are taken from the official releases at HuggingFace(hug, 2024) or the API documentations.
4.4.3. Environment and Model Access
The experiments are performed on a Linux server with 2 Intel®  Xeon®  Platinum 8358P CPUs and 8 NVIDIA A800 GPUs. For open-source LLMs, we deploy a local API server based on vLLM (Kwon et al., 2023), a unified inference and serving engine for LLMs. The weights of open-source models are available at HuggingFace (hug, 2024). To facilitate future research, we list URLs of the models in our replication package. For closed-source GPT models, we invoke the OpenAI API.
4.4.4. Configurations
In Relevant Code Retrieval, we set the maximum chunk size to 2,000 and $k$ (i.e., the number of documents that a retriever should return) to 4 for Java repositories. We set the chunk size to 1,000 and $k$ to 8 for Rust repositories. This is due to the observation that functions and docstrings in our Java repositories are generally longer. In addition, when performing re-ranking, we set the weights of the sparse and dense retrievers to 30% and 70%, respectively.
In evaluations, we set $n$ in pass@$k$ to 10 and $k$ to 1, 3, and 5. To control the randomness of the generated code (Chen et al., 2021), we set the temperature to 0.6 and the nucleus sampling (Holtzman et al., 2019) parameter (top_p) to 0.7. In addition, we set the maximum number of generated tokens to 512.
Figure 5. Two tasks in the Java benchmark, where RepoCoder fails but CatCoder succeeds.
5. Experimental Results
In this section, we analyze the experiment results in detail and answer the research questions.
5.1. RQ1: Effectiveness
5.1.1. Results
Table 3. Evaluation results of different methods on two benchmarks, using CodeLlama-13B-Instruct
Java Rust
pass@k (%) pass@k (%)
Method $k=1$ $k=3$ $k=5$ $k=1$ $k=3$ $k=5$
Vanilla 14.92 22.07 25.21 10.78 15.81 18.15
In-File 35.28 43.77 47.58 41.56 55.70 61.60
RepoCoder 40.95 46.91 49.04 49.44 59.47 63.04
CatCoder 44.67 53.97 57.55 52.67 62.00 65.89
The overall result is presented in Table 3. According to Table 3, we observe that the Vanilla baseline performs poorly as expected because the lack of repository context is likely to cause hallucinations. Next, we can conclude that the In-File baseline performs worse than RepoCoder and CatCoder, as in-file context is not enough to provide key information such as cross-file type structures and API usage. In addition, CatCoder outperforms RepoCoder. In terms of the Java benchmark, the pass@$k$ scores increase by up to 17.35%. For the Rust benchmark, the pass@$k$ scores increase by up to 7.14%.
5.1.2. Qualitative Analysis
To show the advantages of CatCoder over RepoCoder, we also perform a qualitative analysis of the results. Figure 5 presents two cases, in which RepoCoder fails to generate the correct function code while CatCoder succeeds.
Case 1 (method addSerializer): The code generated by RepoCoder compiles but fails to pass unit tests because it uses the wrong field _keySerializers instead of the correct one (i.e., _serializers).
Upon checking the context provided to the LLM, we find that RepoCoder retrieves a similar method named addKeySerializers, which uses the field _keySerializers in its body. The LLM simply copies the code from addKeySerializer and causes the code to fail in unit tests. For CatCoder, it also retrieves the method addKeySerializer, and the pattern of its generated code is the same as the code generated by RepoCoder. However, with the type context extracted by CatCoder, the LLM knows the existence of pairs (field _serializers, method addSerializer) and (field _keySerializers, method addKeySerializers). Therefore, we believe that this makes the LLM filter out some misleading reference code and choose the right field for usage.
Case 2 (method _findWellKnownSimple): The code generated by RepoCoder causes many compilation errors, as it contains non-existent fields and methods in the current scope. In detail, we find that the non-existent CLASS_XXX fields come from another file and RepoCoder retrieves them as context. Without knowing the type context of JavaType and current class (i.e., TypeFactory), the LLM accepts misleading references and generates a non-existent API call stringType().
Answer to RQ1: CatCoder is effective at generating repository-level function code, and it outperforms the RepoCoder baseline by up to 17.35% in terms of pass@$k$ score.
5.2. RQ2: Ablation Study
Table 4. Evaluation results of different variants of our approach on two benchmarks, using CodeLlama-13B-Instruct
Java Rust
pass@k (%) pass@k (%)
Variant $k=1$ $k=3$ $k=5$ $k=1$ $k=3$ $k=5$
FULL 44.67 53.97 57.55 52.67 62.00 65.89
-CR 27.54 38.84 43.66 23.33 35.40 41.04
-TC 39.50 49.48 53.38 49.89 59.61 63.80
To investigate the contribution of the components (as described in Section 3) to the effectiveness of CatCoder, we create the following variants:
  1. (1)
    FULL: The complete version of CatCoder, exactly as Section 3 illustrates.
  2. (2)
    -CR: Relevant Code Retrieval is removed.
  3. (3)
    -TC: Type Context Extraction is removed.
We evaluate the above three variants on the same benchmarks as RQ1, also using CodeLlama-13B-Instruct as the underlying LLM. We compare the incomplete ones with FULL, and the results are illustrated in Table 4. We can conclude that CatCoder benefits from the two components, as the pass@$k$ scores of both incomplete variants drop noticeably. In detail, for the -CR variant, the pass@$k$ scores decrease by up to 38.35% on the Java benchmark and by up to 56.03% on the Rust benchmark. And for the -TC variant, the pass@$k$ scores drop by up to 11.57% on the Java benchmark and by up to 5.3% on the Rust benchmark.
Answer to RQ2: The two components of CatCoder, namely Relevant Code Retrieval and Type Context Extraction, improve the effectiveness of CatCoder.
5.3. RQ3: Generalizability
Figure 6. Evaluation results of the In-File and CatCoder methods on the Java benchmark, using various large language models. The model acronyms on the $x$-axis corresponding to the ones in Table 2. The results are sorted in ascending order based on the values of CatCoder, and the “Average” bar shows the average pass@1 score of all evaluated models.
5.3.1. Results on Generalizability
As described in Section 4.4.2, we select a variety of LLMs for the generalizablity study, and the details of the models are listed in Table 2. Figure 6 shows the pass@$1$ scores of CatCoder on the Java benchmark using the selected LLMs. From the plot, we can notice that CatCoder is able to improve the performance of every model, compared with the In-File context. On average, pass@$1$ scores increase by 32.9%. Therefore, we can conclude that CatCoder can be generalized to different LLMs.
5.3.2. Comparison of different LLMs
We can also learn some facts from pass@$1$ scores across different LLMs [^10] Here, we use the model acronyms in Table 2 to refer to the models.:
❶ Although the selected general purpose LLMs claim to be powerful in many aspects (lla, 2024), they fall behind all code-specialized models except for CG-7B in our benchmarks. One example is ML-8B versus CL-7B. Although CL-7B is fine-tuned from an older foundation model, i.e., Llama 2 (Touvron et al., 2023), it outperforms the new model ML-8B (Llama 3 (lla, 2024)). This indicates that code generation tasks benefit a lot from fine-tuning models using high-quality code datasets.
❷ From the trend in Figure 6, although larger models tend to have a better absolute performance than smaller models, the improvement by CatCoder is more on smaller models. For example, the pass@$1$ scores of CG-7B, ML-8B, CQ-7B, DS-6.7B increase more, compared with SC-15B and CL-34B. This fact provides an insight that we can deploy smaller LLMs on personal devices and use CatCoder to keep up with larger models.
❸ In addition, we can notice that some smaller models even achieve higher performance than larger models. For example, as shown in Figure 6, the performance of two smaller models (i.e., CQ-7B and DS-6.7B) outperforms SC-15B while SC-15B has over twice amount of parameters. This highlights the importance of model architecture and training datasets.
Answer to RQ3: CatCoder is generalizable to many different LLMs, as it improves the pass@$1$ scores of all selected LLMs in our evaluation. In addition, the performance of different LLMs on the benchmark provides valuable insights.
6. Discussion
In this section, we discuss the limitation of our approach CatCoder and the threats to validity.
6.1. When CatCoder fails
Figure 7. A case that CatCoder fails to generate correct code. The first code is generated by CatCoder, and the second one is the ground truth.
Figure 7 shows an example of a failure case generated by CatCoder. The generated code uses all APIs correctly and looks similar to the ground truth code but fails to pass some unit tests because it lacks the logic to handle the null pointers (i.e., the green part in Figure 7). From this case, we can see that function code generation is a complex task, and successful usage of repository context does not always lead to correct code. CatCoder is limited in improving the logical reasoning ability of LLMs, and possible future work could be fine-tuning LLMs with code repositories or using reinforcement learning-based approaches (Le et al., 2022).
6.2. Threats to Validity
6.2.1. Support for More Languages
Our implementation of CatCoder is limited as it only supports Java and Rust. However, the proposed approach does not rely on detailed language specifications and is applicable to any statically typed programming language. In the future, to make CatCoder more practical, we plan to extend it to support more statically typed languages, for example, C++ and C#.
6.2.2. Potential Data Leakage
Our experiments are conducted on the latest state-of-the-art LLMs, and their knowledge cutoff dates are very recent. It is possible that the code repositories in our benchmarks have been included in the training corpora. Thus, the potential data leakage problem could affect the rigor of our results and findings. For example, in this case, it could make the pass@$k$ scores of all approaches higher, as the LLM could output memorized code without having to understand the repository context. In the future, possible improvements may include:
  1. (1)
    Replacing the benchmarks with the most up-to-date ones (Liu et al., 2023; Li et al., 2024) and re-evaluating the approach.
  2. (2)
    Refactoring the benchmarks by renaming the identifiers in the dataset repositories without changing the semantics (Zan et al., 2022).
6.2.3. Efficiency Issue
As CatCoder requires the retrieval and analysis of the source code, there is a noticeable latency from input to output, especially when processing large repositories. Thus, CatCoder is less efficient than commercial tools such as Github Copilot. However, compared to methods that require multiple iterations of model generation, CatCoder’s efficiency is still competitive. In the future, we plan to mitigate this issue by making the components concurrent and caching intermediate data (e.g., embedding vectors and results of type analysis).
7. Related Work
7.1. Large Language Models for Code
Large language models for code (i.e. code LLMs) refer to a family of Transformer (Vaswani et al., 2017) models pre-trained or fine-tuned on massive code datasets. Previous work has shown the exceptional performance of code LLMs in various code-related tasks, such as code generation (Chen et al., 2021), tese case generation (Tufano et al., 2020) and program repair (Xia et al., 2023). In general, pre-trained language models for code can be categorized into the following three types:
  1. (1)
    Encoder-only models, such as CodeBERT (Feng et al., 2020) and GraphCodeBERT (Guo et al., 2020). These models are pre-trained using Masked Language Modeling (MLM) techniques, and are pupular and effective for tasks like classification, where a deep understanding of input data is required.
  2. (2)
    Encoder-Decoder models, such as the CodeT5 family (Wang et al., 2021, 2023). These models first encode the input into an internal representation and then decode this representation into an output sequence.
  3. (3)
    Decoder-only models, such as Codex (Chen et al., 2021), CodeGen (Nijkamp et al., 2022), StarCoder (Li et al., 2023) and CodeLlama (Roziere et al., 2023). These models are pre-trained using Next Token Prediction tasks, and are powerful in generative tasks.
In our experiments, we select a variety of state-of-the-art code LLMs, all of which are decoder-only.
7.2. Benchmarks for Repository-Level Code Generation
Iyer et al. (Iyer et al., 2018) are among the first to highlight the importance of the programmatic context and introduce the task of generating class member functions with natural language descriptions and class contexts. In summary, they construct Concode, a large training dataset consisting of Java classes from online code repositories, and propose an encoder-decoder model for code generation. With the development of large language models, many subsequent repository-level code generation benchmarks arise in academia. Du et al. (Du et al., 2023) make an attempt to evaluate LLMs in the more challenging class-level code generation scenario and manually construct a benchmark named ClassEval, which consists of 100 class-level Python code generation tasks.
Compared with previous benchmarks, for example HumanEval (Chen et al., 2021) and MBPP (Austin et al., 2021), ClassEval is more pragmatic. However, it is still limited because, although the methods in the tasks are not standalone, the programmatic contexts are still restricted to very common libraries and methods within the same file. In real-world scenarios, a code repository generally spans a large number of files, making cross-file contexts essential for LLMs to generate the correct code (Ding et al., 2024).
Toward more realistic evaluation, subsequent benchmarks such as CoderEval (Yu et al., 2024), CrossCodeEval (Ding et al., 2024) and CoderUJB (Zeng et al., 2024) are proposed, all of which are built on a set of real-world open-source repositories and utilize cross-file context for evaluation. A key novelty of CoderEval and CrossCodeEval lies in their ability to support multilingual code generation tasks, while CoderUJB innovatively spans five kinds of practical programming tasks.
As LLMs evolve, Liu et al. (Liu et al., 2023) point out data leakage and memorization problems, which may affect the integrity and trustworthiness of model evaluation results. To mitigate this issue, they propose RepoBench and regularly update the dataset to keep code samples up-to-date. Similarly, Li et al. (Li et al., 2024) propose EvoCodeBench, which includes an automatic pipeline to evolve itself.
7.3. Repository-Level Code Generation with LLMs
Many existing approaches for repository-level code generation perform retrieval. Some work, such as (Zan et al., 2022; Lu et al., 2022), searches for useful information from documentation or external databases. Specifically, Zan et al. (Zan et al., 2022) train a dual-encoder model, which encodes the natural language descriptions and API information, respectively, to retrieve possible APIs from the documentation of the library. Lu et al. (Lu et al., 2022) propose a retrieval-augmented generation approach named ReAcc, which retrieves similar code from a code database they built. However, these approaches are not flexible and general as their retrieval sources (e.g., API documentation and external databases) may not be available in other repositories.
Other retrieval-based approaches explore retrieving similar code snippets within the repositories. Shrivastava et al. (Shrivastava et al., 2023) propose a framework named Repo-Level Prompt Generator to complete the code for holes in single lines. In the framework, they train a prompt proposal classifier that takes the repository files and a set of prompt proposals and outputs the predicted prompt proposal. Zhang et al. (Zhang et al., 2023) propose RepoCoder, which utilizes repository information by iteratively retrieving similar code. In the initial iteration, it uses the natural language descriptions as query input, and in subsequent iterations, it uses previously generated code for query.
Our work CatCoder combines retrieval-based techniques with a new kind of information called type context, and the comparison with RepoCoder in Section 5 shows that type context serves as essential auxillary information for statically type languages. In addition, CatCoder does not require model training or multiple generation iterations, making it more practical in real-world settings.
In addition to the above work, which mainly performs prompt engineering, several model training frameworks (Wu et al., 2024; Guo et al., 2024) and decoding strategies (Agrawal et al., 2024) have been proposed to support repository-level code generation. As our approach CatCoder does not modify the underlying LLM, it can be combined with these frameworks to improve effectiveness.
8. Conclusion and Future Work
In this paper, we aim to address the lack of type dependency information for statically typed programming languages in repository-level code generation. We propose CatCoder, a code generation framework for statically typed programming languages, which improves the performance of LLMs by retrieving relevant code and extracting type context using static analyzers. CatCoder is systematically evaluated on Java and Rust benchmarks that are built from real-world open-source repositories. The results show that CatCoder outperforms the RepoCoder baseline by up to 17.35%. In addition, we evaluate CatCoder with various LLMs. CatCoder improves the performance of all selected LLMs and therefore is generalizable. In future work, possible research attempts include investigating the data leakage problem of the benchmarks and improving the efficiency of CatCoder. These investigations may help further improve the applicability of our proposed approach in practical scenarios.