### v2

# CatCoder: Repository-Level Code Generation with Relevant Code and Type Context

Abstract.
Large language models (LLMs) have demonstrated remarkable capabilities in code generation tasks. However, repository-level code generation presents unique challenges, particularly due to the need to utilize information spread across multiple files within a repository. Specifically, successful generation depends on a solid grasp of both general, context-agnostic knowledge and specific, context-dependent knowledge. While LLMs are widely used for the context-agnostic aspect, existing retrieval-based approaches sometimes fall short as they are limited in obtaining a broader and deeper repository context. In this paper, we present CatCoder, a novel code generation framework designed for statically typed programming languages. CatCoder enhances repository-level code generation by integrating relevant code and type context. Specifically, it leverages static analyzers to extract type dependencies and merges this information with retrieved code to create comprehensive prompts for LLMs. To evaluate the effectiveness of CatCoder, we adapt and construct benchmarks that include 199 Java tasks and 90 Rust tasks. The results show that CatCoder outperforms the RepoCoder baseline by up to 14.44% and 17.35%, in terms of compile@$k$ and pass@$k$ scores. In addition, the generalizability of CatCoder is assessed using various LLMs, including both code-specialized models and general-purpose models. Our findings indicate consistent performance improvements across all models, which underlines the practicality of CatCoder. Furthermore, we evaluate the time consumption of CatCoder in a large open source repository, and the results demonstrate the scalability of CatCoder.
Large Language Model, Code Generation, Repository Context
$c$s: Software and its engineering Automatic programming
1. Introduction
Large language models (LLMs) (wang2021codet5; nijkamp2022codegen; codellama; starcoder2) have been widely applied in various software engineering-related fields, and they have shown remarkable performance in code generation tasks (humaneval; mbpp). Recently, repository-level code generation has attracted much attention (liu2023repobench; wu2024repoformer; zhang2023repocoder; shrivastava2023repository). It refers to the generation of a specific function within a repository and is thus more practical than standalone code generation tasks (i.e., generating functions with no dependencies). This practical scenario introduces unique challenges, as the target code often depends on definitions and usage patterns distributed across multiple files, rather than being confined to a single file. When provided with only the in-file context for such a task, an LLM may hallucinate and produce code that appears plausible but fails to compile or execute correctly within the project (e.g., misusing an API or referencing an undefined identifier). Therefore, a key problem in repository-level code generation is how to provide the LLM with the necessary cross-file context so that it can generate valid and consistent code.
Existing studies have tried to tackle this problem by utilizing repository information with retrieval-based approaches. Zan et al. (private_library) train a dual-encoder model to retrieve API information from library documentation. Lu et al. (lu2022reacc) build an external code database to retrieve similar code. Shrivastava et al.  (shrivastava2023repository) propose RLPG, in which they train a classifier to predict prompt proposals taken from repository files. Zhang et al. (zhang2023repocoder) propose RepoCoder, which iteratively retrieves similar code in a repository. However, these approaches still have limitations. First, methods that depend on external resources or require training specialized retrieval models often lack generalizability across unseen repositories or programming languages. Second, the retrieved code may fail to capture all necessary field accesses and API calls required for correct generation. This issue is particularly pronounced in repositories with minimal code duplication (zhang2023repocoder), or when the desired functionality is novel within the codebase. These limitations underscore the need for more comprehensive modeling of repository context.
In this paper, we aim to address the aforementioned limitations by capturing a broader repository context for the LLM. We observe that, in addition to code retrieval, the types associated with a function can also serve as valuable references for LLMs. Specifically, these related types provide a rich set of fields and methods from which the LLM can select when generating code, thereby enhancing its understanding of repository-specific details. We refer to this explicit information of types as type context. The type context encapsulates localized knowledge embedded in nested scopes, serves as a complement to code retrieval, and helps LLMs develop a deeper understanding of the repository structure. Our work focuses on constructing type context for statically typed programming languages [^1] Statically typed programming languages require variable types to be declared at compile time, such as C, C++, Java, and Rust., where precise type information can be reliably extracted and leveraged. In contrast, dynamically typed languages (e.g., Python and JavaScript) determine variable types at runtime, making type information less accessible and more ambiguous. To extract type context, we utilize existing static analyzers.
Building on the concept of type context, we propose a new repository-level code generation framework named CatCoder. By leveraging the combination of relevant code and type context, CatCoder aims to improve the generation of code that is both logically sound and correctly integrated within its local environment. The framework operates in three stages. First, it retrieves relevant code from the repository. Next, it invokes a static analyzer to obtain dependent types for the target function and constructs a textual representation of the type context. Finally, it integrates both the retrieved code and the type context into a single prompt and generates the target function using an LLM. CatCoder runs on top of a frozen LLM and does not require external databases or additional model training.
To evaluate the effectiveness of CatCoder, we adapt and construct repository-level code generation benchmarks that include 199 Java tasks and 90 Rust tasks in total. We compare CatCoder with vanilla LLM, the In-File context (i.e., contents within the same file), and RepoCoder (zhang2023repocoder), a state-of-the-art framework for repository-level code generation, in terms of compile@$k$ and pass@$k$ (humaneval) scores (i.e., compilation rate and test passing rate). To validate the generalizability of CatCoder, we evaluate CatCoder with various LLMs, including code-specialized models and general-purpose models. To assess the scalability of CatCoder, we conduct extensive evaluations on a large open-source repository, measuring the time consumption of sequential code generation tasks. The results on the benchmarks indicate that CatCoder substantially outperforms the vanilla LLM and the In-File context. In addition, it outperforms RepoCoder by up to 14.44% and 17.35%, in terms of compile@$k$ and pass@$k$. The results of the generalizability study demonstrate that CatCoder improves the performance of all selected LLMs. Furthermore, the results of the scalability study underline the efficiency and practicality of CatCoder in large repositories.
Contributions. In summary, we make the following contributions:
  1. (1)
    We propose a novel framework for statically-typed programming languages, CatCoder, to generate repository-level code by highlighting relevant code and type context in code repositories.
  2. (2)
    To the best of our knowledge, we are the first to construct a repository-level code generation benchmark for Rust, to evaluate the effectiveness of CatCoder on a minor programming language.
  3. (3)
    We evaluate CatCoder using Java and Rust benchmarks, and results indicate that CatCoder outperforms the baselines.
  4. (4)
    We release the replication package [^2] https://github.com/pan2013e/catcoder of CatCoder, including the source code of CatCoder and our benchmarks, to facilitate future research.
Paper Organization. The remainder of the paper is structured as follows. Section 2 introduces the motivation of CatCoder using an example. Section 3 presents the details of CatCoder. Section 4 describes the experimental setups, including baselines, evaluation metrics, benchmarks, and implementation details. Section 5 analyzes the experimental results and provides answers to the research questions. Section 6 discusses the failure cases of CatCoder and threats to validity. Section 7 summarizes the related work. Section 8 concludes the paper with possible future work.
Figure 1. The two essential sources of local context for generating the triu method. (1) Relevant Code (the ones method) provides a crucial, repository-specific pattern for instantiating a RealMatrix object. (2) Type Context provides the equally crucial API definition for the RealMatrix interface, including methods like getEntry needed for element access.
2. Motivating Example
In this section, we present a realistic development scenario to illustrate the core idea behind CatCoder. The motivating example will demonstrate that synthesizing information from both relevant code and type context is crucial for successful generation.
Consider the task of implementing the helper method triu within the CMAESOptimizer class from the widely used Apache Commons Math library, as shown in Fig. 1. As described in its docstring, the triu method is intended to extract the upper triangular part of a given matrix. Its signature is private static RealMatrix triu(final RealMatrix m, int k), where $m$ is the input matrix and $k$ specifies the diagonal offset.
Based on its context-agnostic knowledge of coding, an LLM can infer the general algorithmic principle: the implementation should return a new matrix with the same dimensions, where an element at position $(r,c)$ is copied from $m$ if the column index $c$ is greater than or equal to $r+k$; otherwise, the element is set to zero. However, the requirement of context-dependent knowledge poses challenges to the LLM. The RealMatrix type is an interface in Apache Commons Math and cannot be treated as a nested double array. Proper interaction with a RealMatrix instance requires familiarity with repository-specific API conventions, such as using getEntry(row, column) to access elements. Furthermore, creating a new RealMatrix instance requires using a specific concrete implementation, Array2DRowRealMatrix, and understanding the correct constructor usage. An LLM must therefore grasp these repository-specific details to produce valid code.
Table 1. Different information sources for the triu task
Information Source Key Insight Provided Sufficient Alone?
Info.1 ones method Idiomatic instantiation:
new Array2DRowRealMatrix(...)
gray[1pt/1pt]blackInfo.2 RealMatrix API Correct API access:
m.getRow/ColumnDimension(...)
m.getEntry(...)
gray[1pt/1pt]blackCatCoder (Info.1 + Info.2) Complete, correct implementation
To generate a correct implementation, an LLM must solve two distinct subproblems, that is, (1) how to create the new RealMatrix to be returned, and (2) how to read data from the input RealMatrix m:
  1. (1)
    As shown in Fig. 1, the ones method is retrieved due to its shared return type and location. From this snippet, an LLM can learn a non-trivial pattern for object creation: instantiate a new Array2DRowRealMatrix by first creating a primitive double array and passing it to the constructor. This form of relevant code retrieval reflects a common practice among developers, who often refer to similar methods when writing repository-level code.
  2. (2)
    In addition, Fig. 1 illustrates the available APIs of the RealMatrix type, presented as the type context. Because both the target method triu and the surrounding class CMAESOptimizer reference RealMatrix, its API information provides essential insights. For instance, an LLM can learn how to access dimensions and elements correctly by referencing the provided method signatures. This mirrors another common development strategy: navigating related types (e.g., parameter or return types) to discover usable fields and methods.
Table 2 categorizes the two aforementioned information sources and their key insights. We abstract them as follows:
  1. Info.1
    Relevant code snippets that share similarities in method name, parameters, return types, or docstrings. These examples often perform semantically related or symmetric operations and thus provide reusable implementation patterns.
  2. Info.2
    Type context, defined as the available fields and methods of related types, offering clues for accessing or manipulating object instances correctly.
This motivating example highlights that both Info.1 and Info.2 are indispensable. The absence of either is likely to lead to incorrect or incomplete code generation. While prior work (lu2022reacc; zhang2023repocoder) has leveraged retrieval-augmented generation to utilize relevant code (Info.1), they typically overlook Info.2 (type context). As Fig. 1 shows, retrieved code alone may not fully capture the required API usage. Moreover, in repositories that lack code duplication, relying solely on retrieval is less effective. In such cases, the absence of type context increases the risk of hallucination, such as referencing non-existent methods or fields. To address this gap, CatCoder integrates both Info.1 and Info.2. Our framework is carried out in the following steps:
  1. Step a.
    Relevant Code Retrieval. We retrieve similar code in a repository based on the query (i.e., the docstrings and signature of the function to be completed).
  2. Step b.
    Type Context Extraction. We utilize static analyzers to determine the dependent types and then create type context by extracting the fields and method signatures.
  3. Step c.
    Code Generation via LLM. We merge the results in Step a and Step b into a comprehensive prompt, and invoke an LLM to obtain the generated code.
Figure 2. Overview of our approach
3. Proposed Approach
The framework of CatCoder is presented in Fig. 2. CatCoder takes a source code repository and an empty function (along with its docstrings, if it exists). The output of CatCoder is the completed code of the function.
In addition to Fig. 2, Algorithm 1 presents the pseudocode of the workflow of CatCoder. CatCoder contains two core components: (1) Relevant Code Retrieval and (2) Type Context Extraction. By utilizing these components, CatCoder is able to gather two kinds of referential information for the LLM. After post-processing the raw output of the LLM, CatCoder returns the generated code.
Input: Code repository $R$, Empty function $f$
Output: The completed code of function $f$
1
21ex
3
docs = code_retrieval($R,f$)
/* Section 3.1 */
contexts = type_context($R,f$)
/* Section 3.2 */
prompt = prepare($f$, docs, contexts)
/* Section 3.3 */
4 code = LLM(prompt)
5 return postprocess(code)
Algorithm 1 CatCoder
3.1. Relevant Code Retrieval
3.1.1. Code Splitting
We traverse all the source code files in the repository and split them into smaller chunks. Starting with complete files, we set a maximum chunk size and recursively split larger fragments into smaller ones that fit in the chunk.
As the structure of the source code is often complex, we cannot split it at an arbitrary position. For example, we should not break in the middle of a statement line or some important code elements (e.g., function definitions). Otherwise, the resulting contents would be less meaningful and difficult to understand. Therefore, we split the code only before the following positions (sorted by priority in descending order):
  1. (1)
    Type definitions (e.g., public class Main in Java)
  2. (2)
    Function definitions (e.g., public static void main(String[] args) in Java)
  3. (3)
    Control flow statements (e.g, if, for and while statements in Java)
  4. (4)
[…]
This rule follows the hierarchy of many programming languages and helps protect code structures during splitting. After splitting, we obtain a corpus of code chunks [^3] We call them “documents” later in this subsection. to perform retrieval.
3.1.2. Hybrid Retrieving
In the second step, we build retrievers to search for relevant code. To achieve better search performance, we follow previous research (luan2021sparse) and ensemble two kinds of retrievers: a sparse one and a dense one. The sparse retriever corresponds to keyword matching, and the dense retriever corresponds to semantic search. Each retriever searches and returns top-kk444kk is a hyperparameter, and we describe the settings in Section 4.4.4. relevant documents (i.e., split code chunks), after a query is constructed by concatenating docstrings and the function signature.
❶ Sparse Retrieval
We choose the bag-of-words model for sparse retrieval, and use the BM25 function (robertson1995okapi; trotman2012towards) to rank the set of documents, based on the query terms. Given a document $D$ and a query $Q$ with terms $t_{1},t_{2},\cdots,t_{n}$, the BM25 score can be defined as follows:
BM25(D, Q) = ∑_t ∈Q IDF(t,D) ⋅TF(t,D) ⋅(k1+1)TF(t,D) + k1⋅(1-b+b⋅—D——D—avg) where $\mathrm{IDF}(t,D)$ is the inverse document frequency weight, $\mathrm{TF}(t,D)$ is the term frequency (i.e., the number of appearing times in the document), $|D|$ is the length of the document, and $k_{1}$ and $b$ are hyperparameters. By default, the values of $k_{1}$ and $b$ are set to 1.5 and 0.75. $\mathrm{IDF}(t,D)$ can be calculated by:
IDF(t, D) = logND-DF(t, D) + 0.5DF(t, D) + 0.5, where $N_{D}$ is the number of documents and $\mathrm{DF}(t,D)$ is the document frequency (i.e., the number of documents that contain the term).
In practice, after splitting the code files from a repository, we apply the aforementioned BM25 function for a query and obtain a list of scores. After that, we select documents with the top-$k$ highest scores.
❷ Dense Retrieval
We employ an embedding model for dense retrieval. The model maps natural language sentences and code snippets into dense vector spaces. To select the most relevant documents, our intuitive idea is to find the nearest neighbors around the embedding vector of the query text. Thus, we use Squared Euclidean Distance to measure the proximity between embedding vectors in high-dimensional spaces, and this metric can be calculated by:
d^2(p,q) = ∥p - q ∥^2 = ∑_i=1^n (p_i - q_i)^2, where $n$ is the number of dimensions in the embedding vectors. A lower value of $d^{2}(\mathbf{p},\mathbf{q})$ indicates that $\mathbf{p}$ and $\mathbf{q}$ are more similar.
In practice, we utilize the embedding model and generate vector representations of all documents and query text. Then we calculate $d^{2}(\mathbf{p},\mathbf{q})$ between the query embedding and all document embeddings and select the top-$k$ nearest documents.
3.1.3. Result Ranking
In the third step, we ensemble the two retrievers and re-rank their outputs using the Reciprocal Rank Fusion (RRF) (rrf) algorithm. We assign weights to the retrievers and calculate the weighted RRF score [^5] We assume that a retriever does not output duplicate documents. However, duplicate documents across retrievers are acceptable. for each retrieved document $D$:
RRF(D) = ∑_i=1^n 1(D)⋅wir(D) + 60, where $w_{i}$ is the weight assigned to a certain retriever, $\mathbf{1}(D)$ is an indicator function [^6] The value is 1 if $D$ is one of the outputs of this retriever, otherwise the value is 0., and $r(D)$ is the original rank (if exists) of $D$ in the outputs of this retriever.
After applying the RRF algorithm, we deduplicate the documents and sort them in descending order based on the RRF scores. Now we obtain the results of the Relevant Code Retrieval component.
3.2. Type Context Extraction
Here, we present the type context and the way to extract type contexts from code repositories using static analyzers [^7] The word “type” is an abstract concept and might be different in various languages. For example, it refers to class, interface, and enum in Java; whereas in Rust, it refers to struct and enum..
3.2.1. Type Context
Figure 3. Dependency Graph and Type context of the triu method in class CMAESOptimizer.
To capture the contextual relationships between types within a code repository, Fig. 3 presents a directed dependency graph $G=(V,E)$, in which each vertex $v\in V$ corresponds to a type defined in the repository (e.g., a class or interface). A directed edge $(A,B)\in E$ is created from type $A$ to type $B$, if $A$ references or depends on $B$ (i.e., $B$ occurs in the body of $A$). The dependency graph is able to identify clusters of contextually related types and provide types that an LLM is likely to use and refer to when generating function code.
In Fig. 3, we build the dependency graph for the triu method in class CMAESOptimizer, starting from its own class. As the graph spans, the vertex set and the edge set will eventually converge, and we can obtain the types that are relevant to CMAESOptimizer. To pass this structural information to the LLM, we create type context from the graph by extracting the fields and function signatures within the types and concatenating them into a string. Specifically, for the string representation, we follow the format proposed by Tufano et al. (tufano2020unit), which has been applied in the field of unit test case generation. The right part of Fig. 3 shows the resulting textual type context for the target method (simplified for illustration).
In large repositories, the full dependency graph can become vast, making the resulting type context too long to fit in an LLM’s context window. However, the relevance between types tends to decrease with the distance between them in the graph. For example, in Fig. 3, programmers working on class CMAESOptimizer are more likely to directly use the API of RealMatrix, a direct neighbor, than the API of RealVector, which is an indirect neighbor at a distance of 2. Therefore, in practice, we constrain the scope of type context creation to the direct neighbors of the initial types in the graph. We also prune types from standard libraries, as we assume LLMs possess general knowledge of these libraries from their pre-training.
3.2.2. Building Type Context via Static Analyzer
To get detailed information (i.e., fields and methods) of a type from its name, some straightforward heuristics might be used, e.g., locating imports or looking for similar file names. However, these heuristics are not sound and do not work with all statically typed programming languages. Thus, we need to perform an analysis to accurately navigate to a type and find its definitions and implementations. Many language servers serve as static analyzers and provide this type-based analysis, such as clangd (clangd) for C++, Eclipse JDT (eclipsejdtls) for Java, and rust-analyzer (rust_analyzer) for Rust.
Specifically, to build the type context for a function code generation task, we first determine the set of relevant types (as Section 3.2.1 describes). We start with the signature and form a set of types, and the initial element is the type to which the current function belongs. Then we implement a parser that works collaboratively with the static analyzer. The static analyzer provides detailed type information to the parser. The parser then searches for a new occurring type, adds it to the set, and queries the analyzer with its location (i.e., byte offset or line and column offsets in a file). This interactive process continues until all direct neighbors of the initial set of types have been discovered. Finally, as described in Section 3.2.1, we concatenate the fields and method signatures within the set of types and obtain the type context.
3.3. Code Generation via LLM
Figure 4. Our prompt design. The actual prompt ends right after the signature.
Finally, we describe how to utilize the previous results and invoke the LLM to generate code. Fig. 4 illustrates our prompt design, in which there are two fields: CONTEXT and CODE. To construct a prompt, we follow the context format in previous work (zhang2023repocoder; shrivastava2023repository; liu2023repobench), and fill the CONTEXT field by concatenating the type context and the retrieved code chunks. Then we place the docstrings and the signature of the method at the beginning of the CODE field. The prompt serves as input to a frozen LLM [^8] We discuss the details of selected models and model access in Section 4.4., and we expect the model to complete the CODE field by generating the method body (i.e., the green part in Fig. 4). In addition, to post-process the raw output of the LLM, we employ several procedures [^9] The post-processing procedures may vary with different LLMs and programming languages., including truncating texts, prepending function signatures, and removing Markdown syntax.
4. Experimental Setup
In this section, we first formulate four research questions (RQs). Based on the RQs, we describe the details of the baselines, evaluation metrics, benchmarks, and implementation of our approach.
We aim to answer the following research questions:
  1. RQ1
    How effective is our approach at generating function code in a repository, compared with the baselines? And does our approach generalize to different LLMs?
  2. RQ2
    How do the two components in our approach (i.e., relevant code retrieval and type context extraction) contribute to the overall effectiveness, respectively?
  3. RQ3
    Is our approach generalizable to various large language models? And how do different models perform on the benchmark?
  4. RQ4
    Is our approach efficient when generating a large number of functions in large repositories?
In brief, we evaluate the effectiveness of our approach in RQ1, and then conduct an ablation study in RQ2. In RQ3, we study the generalizability of CatCoder and the performance of different models. Finally, in RQ4, we evaluate the scalability of our approach. For implementation and evaluation of effectiveness, we select two statically typed programming languages, Java and Rust, since Java is one of the most widely used programming languages (tiobe), and Rust is a new system-level language that attracts the attention of many developers (rustlang).
4.1. Baselines
As described in Section 3.3, Fig. 4 shows the prompt template used in our evaluation. As the CONTEXT field is the only variable, we compare our approach with methods that fill different contexts in the prompt. For the repository-level code generation task, there are three types of context: (1) No context, (2) Local context, and (3) Cross-file context. Therefore, we design the following baselines for comparison in RQ1:
  1. (1)
    Vanilla: In this method, we prompt LLMs without any repository context. In this case, all natural language descriptions are given, but the CONTEXT field in Fig. 4 is left empty.
  2. (2)
    In-File: In this method, we consider code generation as a left-to-right file completion task. We extract code before the function to be generated and put it in the CONTEXT field.
  3. (3)
    RepoCoder: RepoCoder (zhang2023repocoder) is an approach that augments the prompt by iteratively retrieving referential code fragments, and it is originally proposed for Python code repositories. We adopt their idea and implement this baseline for Java and Rust based on our retrieval pipeline. Specifically, the initial iteration uses natural language descriptions (i.e., docstrings) as the query for code retrieval. From the second iteration, it uses the previously generated code for retrieval.
4.2. Evaluation Metrics
For the function code generation task, we argue that metrics based on text similarity are not accurate, and it is essential to use metrics based on compilation and execution results. Therefore, we use compilation rate and test passing rate, which evaluate syntax correctness and functional correctness, respectively. Specifically, to calculate such a rate, we follow Chen et al. (humaneval) and use the unbiased score@$k$. It is defined as
score@$k$ = E_Tasks [ 1 - (n-ck)(nk) ], where $n$ is the number of code samples generated per task and $k$ is the parameter in “score@kk” ($n\geq k$). Based on this general formula, we define two detailed metrics that correspond to compilation rate and test passing rate:
  • compile@k. In the formula, $c$ refers to the number of syntactically correct code samples. A code sample is syntactically correct if and only if it passes compiler checks and compiles.
  • pass@k. In the formula, $c$ refers to the number of functionally correct code samples. A code sample is functionally correct if and only if it passes the tests in the benchmark.
We calculate score@$k$ values for each task in the benchmark, and the final score@$k$ for a benchmark is the mean value of all score@$k$ values.
4.3. Benchmarks
4.3.1. Java Benchmark
There are several publicly available repository-level benchmarks for Java (ding2024crosscodeeval; yu2024codereval; agrawal2024monitor; zeng2024coderujb). However, it is challenging to use many of them to evaluate CatCoder because they either lack the raw repository code or are not designed for function code generation tasks. Considering the challenges, the CoderUJB (zeng2024coderujb) benchmark is a suitable base benchmark that contains 238 code generation tasks, and we successfully adapt it for our evaluation. In the adaptation process, we filter out a few functions that satisfy one of the following patterns:
  1. (1)
    The length of function code exceeds the max output limit (i.e., 512 tokens in our settings).
  2. (2)
    The function does not have any modifiers or annotations. This helps us to perform post-processing more easily.
  3. (3)
    The function is standalone, which means that it can be moved to a new class [^10] We assume that all Java standard libraries are imported in advance when applying this rule. and still compiles. The removal of standalone functions makes the benchmark more complex and difficult, and helps better evaluate the performance of repository-level code generation.
We apply the filtering rules sequentially: 17 functions are filtered out by the first rule, 5 by the second, and 17 by the third. As a result, after removing these 39 functions, we obtain 199 coding tasks. As the methods in the benchmark are mined by the CoderUJB authors from the Defects4J (just2014defects4j) dataset, it is convenient to evaluate the correctness of the generated code using the existing extensive unit test methods. Specifically, we “patch” the repository with generated code and run the defects4j test -t ... command.
4.3.2. Rust Benchmark
To our knowledge, the only code generation benchmark for the Rust programming language is MultiPL-E (cassano2023multipl), which translates the original HumanEval (humaneval) and MBPP (mbpp) benchmarks. However, this benchmark is designed for standalone function code generation tasks and thus is not suitable for our evaluation. In this case, we construct a new benchmark named RustEval.
Table 2. The crates used in our Rust benchmark
Crate # Downloads Crate # Downloads
bit-vec 54M sdl2 1.7M
curv-kzen 161K splay_tree 72K
dlv-list 14M sshkeys 189K
gptman 514K unix_path 86K
mathru 202K vec_map 81M
multiset 99K without-alloc 259K
rocket_http 4.2M
As shown in Table 2, we collect 13 repositories of various categories from crates.io (cratesio), the official Rust package registry. In these repositories, we obtain 90 functions with docstrings that satisfy the official guidelines (rustdoc) (i.e., have detailed descriptions and at least one documentation test). To test the correctness of the generated code, we utilize the documentation tests in the docstrings by running the cargo test --doc ... command.
4.3.3. Benchmark Statistics
Table 3. Statistics of our benchmarks, including number of tasks, average number of lines of code (NLOC) per task, and average cyclomatic complexity number (CCN) per task.
Java Rust
# Tasks 199 90
Avg. NLOC 16.59 10.68
Avg. CCN 4.59 2.64
Table 3 presents the statistics of our benchmarks, where we assess their complexity in terms of the average number of lines of code (NLOC) and cyclomatic complexity number (CCN) per task. Compared to the code completion task in our baseline paper, which requires generating only a single line of code, our function code generation task involves producing more lines and deeper code structures, as reflected in the NLOC and CCN values shown in Table 3. In addition, compared to the Java benchmark, the Rust benchmark is less syntactically verbose and structurally simpler. We attribute this to idiomatic Rust’s reduced boilerplate and more concise abstractions. However, from a semantic standpoint, this does not imply that code generation is easier, as the underlying logic can still be dense and complex.
4.4. Implementation
4.4.1. Approach
We implement CatCoder for Java and Rust, and the source code is written in Python and Rust. Specifically in Relevant Code Retrieval, we use MPNet V2 (mpnetv2) as the embedding model. It is an open-source model fine-tuned from the base MPNet (song2020mpnet) model using massive training data, including the popular CodeSearchNet (husain2019codesearchnet). For evaluation on the benchmarks, the ground truth code in the repositories is removed during retrieval to prevent leakage. In Type Context Extraction, we use Eclipse JDT.LS (eclipsejdtls) and rust-analyzer (rust_analyzer) as the static analyzers. For Java, our approach interacts with Eclipse JDT.LS through the Language Server Protocol (lsp_spec). For Rust, we use rust-analyzer as a third-party library and implement a binding between Rust and Python.
4.4.2. Model Selection
Table 4. Models used for evaluation in RQ3. The model acronyms correspond to the $x$-axis labels in Figure 6.
Model ID Acronym Open-Source?
Code Specialized
(codellama) CodeLlama-7b-Instruct CL-7B
(codellama) CodeLlama-13b-Instruct CL-13B
(deepseekcoder) deepseek-coder-6.7b-instruct DS-6.7B
(codegemma) codegemma-7b-it CG-7B
(codeqwen) CodeQwen1.5-7B-Chat CQ-7B
General Purpose
(llama3) Meta-Llama-3-8B-Instruct ML-8B
For RQ1 and RQ2, we perform the evaluation using CodeLlama-13B-Instruct (codellama) as the default LLM, due to its state-of-the-art performance among models of medium size. For RQ3, to evaluate the generalizability of CatCoder, we select a variety of state-of-the-art LLMs with different sizes, including code-specialized models and general-purpose models. The details of the models are listed in Table 4, in which the model IDs are taken from the official releases at HuggingFace(huggingface) or the API documentations.
4.4.3. Environment and Model Access
The experiments are performed on a Linux server with capable NVIDIA GPUs. For open-source LLMs, we deploy a local API server based on vLLM (vllm), a unified inference and serving engine for LLMs. The weights of open-source models are available at HuggingFace (huggingface). To facilitate future research, we list URLs of the models in our replication package.
4.4.4. Configurations
In Relevant Code Retrieval, we set the maximum chunk size to 2,000 and $k$ (i.e., the number of documents that a retriever should return) to 4 for Java repositories. We set the chunk size to 1,000 and $k$ to 8 for Rust repositories. This is due to the observation that functions and docstrings in our Java repositories are generally longer. In addition, when performing re-ranking, we set the weights of the sparse and dense retrievers to 30% and 70%, respectively.
In evaluations, we set $n$ in compile@$k$ and pass@$k$ to 10 and $k$ to 1, 3, and 5. To control the randomness of the generated code (humaneval), we set the temperature to 0.6 and the nucleus sampling (nucleus) parameter (top_p) to 0.7. In addition, we set the maximum number of generated tokens to 512.
5. Experimental Results
In this section, we analyze the experiment results in detail and answer the research questions.
5.1. RQ1: Effectiveness
5.1.1. Results
Table 5. Evaluation results of different methods on two benchmarks, using CodeLlama-13B-Instruct
Java Rust
compile@k (%) pass@k (%) compile@k (%) pass@k (%)
Method 11 33 55 11 33 55 11 33 55 11 33 55
Vanilla 33.6 49.4 56.2 14.9 22.1 25.2 18.0 28.3 33.1 10.8 15.8 18.2
In-File 59.7 72.7 76.8 35.3 43.8 47.6 52.1 66.5 70.9 41.6 55.7 61.6
RepoCoder 63.0 74.1 78.2 41.0 46.9 49.0 63.7 73.4 76.4 49.4 59.5 63.0
CatCoder 71.9 84.8 88.4 44.7 54.0 57.5 65.1 75.5 78.9 52.7 62.0 65.9
The overall result is presented in Table 5. According to Table 5, we observe that the Vanilla baseline performs poorly as expected because the lack of repository context is likely to cause hallucinations. Next, we can conclude that the In-File baseline performs worse than RepoCoder and CatCoder, as in-file context is not enough to provide key information such as cross-file type structures and API usage. In addition, CatCoder outperforms RepoCoder. In terms of the Java benchmark, the compile@$k$ and pass@$k$ scores increase by up to 14.44% and 17.35%, respectively. For the Rust benchmark, the compile@$k$ and pass@$k$ scores increase by up to 3.27% and 6.68%.
Figure 5. Two tasks in the Java benchmark, where RepoCoder fails but CatCoder succeeds.
5.1.2. Qualitative Analysis
To show the advantages of CatCoder over RepoCoder, we also perform a qualitative analysis of the results. Fig. 5 presents two cases, in which RepoCoder fails to generate the correct function code while CatCoder succeeds.
Case 1 (method addSerializer): The code generated by RepoCoder compiles but fails to pass unit tests because it uses the wrong field _keySerializers instead of the correct one (i.e., _serializers).
Upon checking the context provided to the LLM, we find that RepoCoder retrieves a similar method named addKeySerializers, which uses the field _keySerializers in its body. The LLM simply copies the code from addKeySerializer and causes the code to fail in unit tests. For CatCoder, it also retrieves the method addKeySerializer, and the pattern of its generated code is the same as the code generated by RepoCoder. However, with the type context extracted by CatCoder, the LLM knows the existence of pairs (field _serializers, method addSerializer) and (field _keySerializers, method addKeySerializers). Therefore, we believe that this makes the LLM filter out some misleading reference code and choose the right field for usage.
Case 2 (method _findWellKnownSimple): The code generated by RepoCoder causes many compilation errors, as it contains non-existent fields and methods in the current scope.
In detail, we find that the non-existent CLASS_XXX fields come from another file, and RepoCoder retrieves them as context. Without knowing the type context of JavaType and current class (i.e., TypeFactory), the LLM accepts misleading references and generates a non-existent API call stringType().
5.1.3. Comparison of Java and Rust
As shown in Table 5, the magnitude of improvement provided by CatCoder over the RepoCoder baseline differs between the two languages. On the Java benchmark, CatCoder achieves a relative improvement of up to 17.35% in pass@k, whereas on Rust, the maximum relative improvement is 6.68%. This suggests that CatCoder offers a more substantial advantage in the context of the evaluated Java projects. We attribute this difference to the characteristics of the languages and the structural properties of the benchmark projects. As indicated in Table 3, the Java benchmark is, on average, more structurally complex and verbose than the Rust benchmark. We argue that this increased complexity creates scenarios in which the type context provided by CatCoder becomes a more critical factor for successful code generation, as illustrated in Section 5.1.2.
Table 6. Results of CodeBLEU, using CodeLlama-13B-Instruct. C: CodeBLEU score; N: n-gram match score; W: weighted n-gram match score; S: syntax match score; D: dataflow match score.
Java Rust
Method C N W S D C N W S
Vanilla 0.36 0.23 0.30 0.41 0.37 0.25 0.16 0.21 0.39
In-File 0.51 0.40 0.46 0.58 0.55 0.38 0.29 0.33 0.54
RepoCoder 0.57 0.47 0.53 0.64 0.61 0.46 0.37 0.42 0.60
CatCoder 0.56 0.46 0.52 0.63 0.59 0.44 0.35 0.40 0.60
5.1.4. Semantic Similarity vs. Test Pass Rate
In addition to compile@k and pass@k, we conduct an evaluation using CodeBLEU (ren2020codebleu), a metric that lies between purely text similarity-based and execution-based metrics. Beyond traditional n-gram matching, CodeBLEU incorporates syntactic and semantic matching, thus better capturing the semantic similarity between code snippets. It is computed as: CodeBLEU = α⋅BLEU + β⋅BLEU_weight + γ⋅Match_ast + δ⋅Match_df, where $\mathrm{BLEU}$ is the n-gram match score, $\mathrm{BLEU_{weight}}$ is the weighted n-gram match score, $\mathrm{Match_{ast}}$ is the abstract syntax tree (AST) match score, and $\mathrm{Match_{df}}$ is the dataflow match score. The parameters $\alpha,\beta,\gamma,\text{and }\delta$ are weighting coefficients.
We use an open-source implementation [^11] https://github.com/k4black/codebleu of CodeBLEU, and the results are presented in Table 6. For the Java benchmark, the weights are set to $\alpha=\beta=\gamma=\delta=0.25$. For the Rust benchmark, since the tool cannot extract dataflow from Rust code, we omit the dataflow match component and assign the weights to $\alpha=\beta=\gamma=0.33$ and $\delta=0$.
On both Java and Rust benchmarks, CatCoder and RepoCoder outperform the other two baselines, Vanilla and In-File, across CodeBLEU and all four sub-metrics, suggesting that relevant code retrieval helps LLMs generate code more similar to the ground truth. Notably, the scores of CatCoder and RepoCoder are very close, with RepoCoder even slightly outperforming CatCoder in CodeBLEU. However, since CatCoder achieves higher compile@k and pass@k scores, particularly on the Java benchmark, this indicates that RepoCoder sometimes produces code structurally similar to the ground truth but fails due to subtle issues. This underscores the importance of execution-based metrics in evaluating repository-level code generation.
Answer to RQ1: CatCoder is effective at generating repository-level function code, and it outperforms the RepoCoder baseline by up to 14.44% and 17.35% in terms of compile@$k$ and pass@$k$ scores. Besides, in qualitative analysis, we find that the type context is effective in filtering out misleading reference code.
5.2. RQ2: Ablation Study
Table 7. Evaluation results of different variants on two benchmarks, using CodeLlama-13B-Instruct
Variant Java Rust
compile@k (%) pass@k (%) compile@k (%) pass@k (%)
11 33 55 11 33 55 11 33 55 11 33 55
FULL 71.9 84.8 88.3 44.7 54.0 57.5 65.1 75.5 78.9 52.7 62.0 65.9
-CR 57.9 75.4 81.4 27.5 38.8 43.7 35.0 50.7 56.3 23.3 35.4 41.0
-TC 67.1 80.3 84.8 39.5 49.5 53.4 59.6 70.1 73.8 49.9 59.6 63.8
To investigate the contribution of the components (as described in Section 3) to the effectiveness of CatCoder, we create the following variants:
  1. (1)
    FULL: The complete version of CatCoder, exactly as Section 3 illustrates.
  2. (2)
    -CR: Relevant Code Retrieval is removed.
  3. (3)
    -TC: Type Context Extraction is removed.
We evaluate the above three variants on the same benchmarks as RQ1, also using CodeLlama-13B-Instruct as the underlying LLM. We compare the incomplete ones with FULL, and the results are illustrated in Table 7. We can conclude that CatCoder benefits from the two components, as the compile@$k$ and pass@$k$ scores of both incomplete variants drop noticeably. In detail, for the -CR variant, the compile@$k$ and pass@$k$ scores decrease by up to 19.43% and 38.35%, respectively, on the Java benchmark, and by up to 46.24% and 56.03% on the Rust benchmark. And for the -TC variant, the compile@$k$ and pass@$k$ scores drop by up to 6.63% and 11.57% respectively on the Java benchmark, and by up to 8.52% and 5.3% on the Rust benchmark.
Notably, removing Type Context Extraction leads to a smaller performance decline compared to removing Relevant Code Retrieval. This observation underscores the effectiveness of our hybrid retrieval mechanism. However, we emphasize that the importance of type context is not diminished. In many instances, particularly those involving functionalities that are symmetric to, or minor variations of, existing code in the repository, the retriever is able to find highly relevant examples. The presence of duplicated code snippets benefits retrieval-based generation by providing the LLM with strong, nearly complete templates for the target function. In such cases, the retrieved code is sufficient for the LLM to produce a correct solution, rendering the additional type context less critical. However, the performance gap between the -TC variant and FULL highlights a subset of more challenging problems where retrieval alone is insufficient. The qualitative examples in Fig. 5 illustrate a typical failure mode: the retrieved code may include distracting snippets or reference elements outside the current scope, resulting in generation errors. Therefore, we argue that our Type Context Extraction component is an essential and complementary mechanism.
Answer to RQ2: The two components of CatCoder, namely Relevant Code Retrieval and Type Context Extraction, improve the effectiveness of CatCoder.
5.3. RQ3: Generalizability
Figure 6. pass@1 performance comparison using various LLMs on the Java and Rust benchmarks. Subfigures (a) and (b) compare CatCoder with baseline methods—In-File and RepoCoder—while (c) and (d) compare CatCoder with its variants. Model acronyms on the $x$-axis are consistent with those listed in Table 4, and variant names refer to those introduced in RQ2. Models are sorted in ascending order based on the performance of CatCoder. The “Average” bar represents the mean pass@1 score across all evaluated models.
To assess whether the benefits of CatCoder are specific to a single LLM or generalize across different models, we conduct an extensive generalizability study. As described in Section 4.4.2, we select a diverse set of LLMs, including both code-specialized and general-purpose models of varying sizes (Table 4). We then evaluate the performance of the In-File, RepoCoder, and CatCoder approaches, as well as the CatCoder variants (-CR and -TC), using each of these models on both our Java and Rust benchmarks.
5.3.1. Results on Generalizability
Fig. 6 shows the pass@$1$ scores for these experiments. The results for both Java (Fig. 6a, 6c) and Rust (Fig. 6b, 6d) benchmarks suggest that CatCoder is able to improve the performance of every tested model.
On the Java benchmark, CatCoder improves the average pass@$1$ score by 13.73%, compared to RepoCoder (Fig. 6a). The improvement is universal, ranging from a significant increase for models like DS-6.7B to a more modest increase for models like CL-13B. Similarly, on the Rust benchmark (Fig. 6b), CatCoder consistently outperforms both In-File and RepoCoder for all LLMs. It improves the average pass@$1$ score by 8.4% compared to RepoCoder. Additionally, the relative performance ranking of the models remains largely consistent across both languages, with CL-13B and DS-6.7B generally at the top, and with CG-7B and ML-8B showing more modest results.
Furthermore, the ablation study results, when generalized across models (Fig. 6c, 6d), are consistent with the findings from RQ2. For both Java and Rust, the -CR variant (without Code Retrieval) consistently leads to the most significant performance degradation, while the -TC variant (without Type Context) shows a smaller but still noticeable drop. On average, across all models on the Java benchmark, the pass@$1$ score decreases by 46.89% without Relevant Code Retrieval, and decreases by 17.17% without Type Context Extraction. This pattern holds for Rust as well. On average, removing Relevant Code Retrieval decreases the pass@1 score by 57.94%, while removing Type Context Extraction leads to an 8.21% decrease.
In conclusion, our results support the generalizability of CatCoder. It consistently enhances the performance of a wide variety of LLMs on both Java and Rust, demonstrating that the architectural benefits of CatCoder are model-agnostic.
5.3.2. Comparison of different LLMs
The results of our generalizability study also offer insights into the current landscape of LLMs for code:
❶ Although the selected general purpose LLM (i.e., ML-8B) claims to be powerful in many aspects (llama3), it falls behind all selected code-specialized models except for CG-7B in our benchmarks. One example is ML-8B versus CL-7B. Although CL-7B is fine-tuned from an older foundation model (i.e., Llama 2 (touvron2023llama2)), it outperforms the new model ML-8B (i.e., Llama 3 (llama3)). This indicates that code generation tasks benefit a lot from fine-tuning models using high-quality code datasets.
❷ From the trend in Fig. 6, although larger models tend to have a better absolute performance than smaller models, the improvement by CatCoder is larger on smaller models. For instance, the pass@1 score increase for models like ML-8B, CQ-7B, and DS-6.7B is more pronounced than the increase for CL-13B. This suggests that providing high-quality, structured context via frameworks like CatCoder can help smaller or less specialized models bridge the performance gap with larger, more powerful ones.
Answer to RQ3: CatCoder is generalizable to many different LLMs, as it improves the pass@$1$ scores of all selected LLMs, compared with both In-File method and RepoCoder. In addition, the performance of different LLMs on the benchmark provides valuable insights.
5.4. RQ4: Scalability
Figure 7. Detailed latency of the components when constructing prompts for 29 code generation tasks sequentially in the closure-compiler project.
Table 8. Average latency of the components when constructing prompts for 29 code generation tasks sequentially in the closure-compiler project. CR: Relevant Code Retrieval, in Section 3.1; TC: Type Context Extraction, in Section 3.2.
Case CR TC Total
Cold start (Task 0) 38.8s 4.62s 38.8s
Avg. after cold start 1.08s 1.32s 1.56s
Avg. speedup w/ cache 36x 3.5x 25x
In real-world software development, the latency of repository-level code generation tools is crucial, as they inherently process considerable volumes of contextual code in large and complex projects. Significant latencies would negate the advantages of automated assistance and lead to user frustration. Therefore, we propose this research question and investigate the scalability of CatCoder when applied to functions within large-scale code repositories.
5.4.1. Experimental Design
To robustly assess CatCoder’s scalability, the closure-compiler repository [^12] The Closure Compiler is a tool for checking and optimizing JavaScript code. The project is developed by Google and has attracted wide attention. in our benchmark (Section 4.3) is selected for evaluation, by considering the following aspects:
  • This repository is the largest one in our benchmark and comprises approximately 400,000 lines of code, representing a non-trivial, real-world system. Thus, it can present a genuine scalability challenge for CatCoder.
  • It is a well-known and actively developed open-source project. Thus, the evaluation is realistic and can reflect the types of complex dependencies, code structures, and API usage patterns that CatCoder would encounter in practical deployment.
Specifically, our benchmark contains 29 distinct function code generation tasks for the closure-compiler repository. For evaluation, we open this repository, invoke CatCoder on these code generation tasks sequentially, and examine the time consumption of each code generation task. This sequential execution simulates a common developer workflow where a programmer might work on several functions or address multiple issues within the same project during a single coding session.
For time consumption, we focus on the latency of CatCoder’s core components (i.e., Relevant Code Retrieval and Type Context Extraction), while excluding the inference latency of the underlying LLM. This is due to that the time taken for LLM inference is highly dependent on external factors that are largely independent of CatCoder, including the specific LLM chosen, the underlying hardware, and the deployment environment. In addition, the length of prompts constructed by CatCoder does not scale with repository sizes, making the inference latency less important in scalability evaluation.
5.4.2. Analysis of Scalability
Fig. 7 shows the detailed latency of the two components of CatCoder when running each sequential code generation task in the closure-compiler repository. In addition, Table 8 shows the statistics of average latency. It is also worth noting that the two components are independent and can run in parallel. Thus, the total time is bounded by the slower component.
We can notice that CatCoder takes a longer time to process the very first code generation task in a repository that it has not previously analyzed. We refer to the first task as “cold start”. The latency of cold start can be attributed to several setup operations in our approach:
  • For Relevant Code Retrieval (CR), the setup operations mainly involve generating embedding vectors for code chunks, which are crucial for subsequent semantic search.
  • For Type Context Extraction (TC), the operations include the static analyzer’s setup processes, such as parsing the codebase and building representations of type hierarchies and dependencies.
As shown in Table 8, the measured cold start time is 38.8 seconds, with the Relevant Code Retrieval component as the main bottleneck. Considering that the closure-compiler repository is large, and it also takes many IDEs to index a new large repository for a long time, we believe that the cold start overhead of CatCoder is still acceptable for enabling its advanced context-aware code generation capabilities.
Following the initial cold start, CatCoder’s performance for subsequent tasks improves significantly. Table 8 shows the reduction in average latency for tasks after the first one: Relevant Code Retrieval time drops to 1.08s, Type Context Extraction to 1.32s, and the total prompt construction time averages 1.56s. The results indicate an average 36x speedup for Relevant Code Retrieval, 3.5x speedup for Type Context Extraction, and 25x speedup for total time. In this case, after the initial waiting, users do not have to wait for a long time each time to receive a response in an interactive development workflow. This is due to CatCoder’s caching mechanism, which stores intermediate data, including embedding vectors of code chunks and intermediate states of the static analyzer. The cached data is computationally expensive to generate, but generally stable across a sequence of related development tasks within a single coding session. The massive speedup after the cold start demonstrates the effective reuse of cache data.
Answer to RQ4: CatCoder shows practical scalability at a series of code generation tasks within large repositories. In the case of the closure-compiler repository, the caching mechanism of CatCoder is able to introduce a 25x average speedup after the cold start.
6. Discussion
In this section, we discuss the limitations of our approach and outline potential directions for future work, followed by a discussion of threats to validity.
6.1. Limitations of CatCoder
6.1.1. Limitation in the Underlying LLM
Figure 8. A case that CatCoder fails to generate correct code. The first code is generated by CatCoder, and the second one is the ground truth.
Apart from relevant code and type context, an LLM’s internal capability of understanding general syntax and algorithmic principles is equally essential for correct repository-level code generation. Similar to many prior studies (liu2023repobench; wu2024repoformer; zhang2023repocoder; shrivastava2023repository), CatCoder operates on a frozen LLM. As such, this line of work is inherently limited in enhancing the model’s logical reasoning abilities.
Fig. 8 presents a failure case produced by CatCoder. Although the generated code correctly uses all relevant APIs and closely resembles the ground truth implementation, it fails to pass some unit tests due to missing the logic for handling null pointers (highlighted in green in Fig. 8). This example illustrates that function code generation is inherently complex, and the effective use of repository context does not always guarantee correctness. A possible direction for future work could be fine-tuning LLMs on code repositories to enhance repository-level in-context learning, or adopting reinforcement learning approaches (le2022coderl) to improve reasoning and coding capabilities.
6.1.2. Limitation in Retrieval
In the Relevant Code Retrieval component, we adopt a hybrid retrieval strategy that combines sparse (BM25) and dense (embedding-based) methods to achieve high recall and precision. However, like any information retrieval system, it is not infallible. The retriever may still return incomplete or even misleading context. To some extent, the type context in CatCoder helps mitigate this limitation by supplying accessible API information that the LLM can cross-reference, as illustrated in Section 5.1.2. Possible future work includes enhancing the retrieval and reranking pipeline using powerful LLMs.
6.1.3. Limitation in Static Analysis
Although Java is a statically typed programming language, it supports dynamic features such as reflection. The Java Reflection API enables dynamic class loading, object instantiation, and method invocation based on runtime strings. Despite the capabilities of the static analyzer we employ (i.e., Eclipse JDT.LS, commonly used in IDEs), handling these dynamic features remains challenging. For example, if a target function is related to a field whose class is loaded reflectively at runtime (e.g., via the Class.forName method), the analyzer may fail to resolve the actual class and its members. In such cases, the resulting type context may omit information related to the truly dependent class. Therefore, a limitation is that CatCoder may produce incomplete type context when the repository heavily relies on dynamic features.
6.2. Threats to Validity
6.2.1. Internal Threats
In general, static analyzers may introduce unsound results, which can, in turn, affect dependent approaches. However, we argue that this is not a concern in CatCoder. Specifically, we use static analysis only to locate types and obtain their definitions and implementations. This form of analysis is lightweight and significantly more accurate than other forms, such as pointer or reflection analysis, as it operates solely on explicit type names in the code. Besides, the tools we use (i.e., Eclipse JDT and rust-analyzer) and the analysis we apply have been integrated into widely used IDEs and code editors (e.g., Eclipse and Visual Studio Code). Therefore, we consider further verification of the analysis results in CatCoder unnecessary, and we trust that our evaluation results remain valid.
6.2.2. External Threats
Our implementation of CatCoder is limited as it only supports Java and Rust. However, the proposed approach does not rely on detailed language specifications and is applicable to any statically typed programming language. In the future, to make CatCoder more practical, we plan to extend it to support more statically typed languages, for example, C++ and C#.
7. Related Work
7.1. Large Language Models for Code
Large language models for code (i.e., code LLMs) refer to a family of Transformer (vaswani2017attention) models pre-trained or fine-tuned on massive code datasets. Previous work has shown the exceptional performance of code LLMs in various code-related tasks, such as code generation (humaneval), test case generation (tufano2020unit), and program repair (program_repair). In general, pre-trained language models for code can be categorized into the following three types:
  1. (1)
    Encoder-only models, such as CodeBERT (feng2020codebert) and GraphCodeBERT (guo2020graphcodebert). These models are pre-trained using Masked Language Modeling (MLM) techniques, and are popular and effective for tasks like classification, where a deep understanding of input data is required.
  2. (2)
    Encoder-Decoder models, such as the CodeT5 family (wang2021codet5; wang2023codet5+). These models first encode the input into an internal representation and then decode it into an output sequence.
  3. (3)
    Decoder-only models, such as Codex (humaneval), CodeGen (nijkamp2022codegen), StarCoder (li2023starcoder) and CodeLlama (codellama). These models are pre-trained using Next Token Prediction tasks and are powerful in generative tasks.
In our experiments, we select a variety of state-of-the-art code LLMs, all of which are decoder-only.
7.2. Benchmarks for Repository-Level Code Generation
Iyer et al. (concode) are among the first to highlight the importance of the programmatic context and introduce the task of generating class member functions with natural language descriptions and class contexts. In summary, they construct Concode, a large training dataset consisting of Java classes from online code repositories, and propose an encoder-decoder model for code generation. With the development of large language models, many subsequent repository-level code generation benchmarks have arisen in academia. Du et al. (du2023classeval) make an attempt to evaluate LLMs in the more challenging class-level code generation scenario and manually construct a benchmark named ClassEval, which consists of 100 class-level Python code generation tasks.
Compared with previous benchmarks, for example HumanEval (humaneval) and MBPP (mbpp), ClassEval is more pragmatic. However, it is still limited because, although the methods in the tasks are not standalone, the programmatic contexts are still restricted to very common libraries and methods within the same file. In real-world scenarios, a code repository generally spans a large number of files, making cross-file contexts essential for LLMs to generate the correct code (ding2024crosscodeeval).
Toward more realistic evaluation, subsequent benchmarks such as CoderEval (yu2024codereval), CrossCodeEval (ding2024crosscodeeval), and CoderUJB (zeng2024coderujb) are proposed, all of which are built on a set of real-world open-source repositories and utilize cross-file context for evaluation. A key novelty of CoderEval and CrossCodeEval lies in their ability to support multilingual code generation tasks, while CoderUJB innovatively spans five kinds of practical programming tasks.
As LLMs evolve, Liu et al. (liu2023repobench) point out data leakage and memorization problems, which may affect the integrity and trustworthiness of model evaluation results. To mitigate this issue, they propose RepoBench and regularly update the dataset to keep code samples up-to-date. Similarly, Li et al. (li2024evocodebench) propose EvoCodeBench, which includes an automatic pipeline to evolve itself.
7.3. Repository-Level Code Generation with LLMs
Many existing approaches for repository-level code generation perform retrieval. Some work, such as (private_library; lu2022reacc), searches for useful information from documentation or external databases. Specifically, Zan et al. (private_library) train a dual-encoder model, which encodes the natural language descriptions and API information, respectively, to retrieve possible APIs from the documentation of the library. Lu et al. (lu2022reacc) propose a retrieval-augmented generation approach named ReAcc, which retrieves similar code from a code database they built. However, these approaches are not flexible and general as their retrieval sources (e.g., API documentation and external databases) may not be available in other repositories.
Other retrieval-based approaches explore retrieving similar code snippets within the repositories. Shrivastava et al. (shrivastava2023repository) propose a framework named Repo-Level Prompt Generator to complete the code for holes in single lines. In the framework, they train a prompt proposal classifier that takes the repository files and a set of prompt proposals and outputs the predicted prompt proposal. Zhang et al. (zhang2023repocoder) propose RepoCoder, which utilizes repository information by iteratively retrieving similar code. In the initial iteration, it uses the natural language descriptions as query input, and in subsequent iterations, it uses previously generated code for queries.
Our work CatCoder combines retrieval-based techniques with a new kind of information called type context, and the comparison with RepoCoder in Section 5 shows that type context serves as essential auxiliary information for statically typed languages. In addition, CatCoder does not require model training or multiple generation iterations, making it more practical in real-world settings.
In addition to the above work, which primarily focuses on prompt engineering, several model training frameworks (wu2024repoformer; deepseekcoder) and decoding strategies (agrawal2024monitor) have been proposed to support repository-level code generation. As our approach CatCoder does not modify the underlying LLM, it can be combined with these frameworks to improve effectiveness.
8. Conclusion and Future Work
In this paper, we aim to address the lack of type dependency information in repository-level code generation for statically typed programming languages. We propose CatCoder, a code generation framework for statically typed programming languages, which improves the performance of LLMs by retrieving relevant code and extracting type context using static analyzers. CatCoder is systematically evaluated on Java and Rust benchmarks that are built from real-world open-source repositories. The results show that CatCoder outperforms the RepoCoder baseline by up to 17.35% in terms of pass@k scores. Further studies demonstrate the generalizability and scalability of CatCoder. We release the replication package of CatCoder, including the source code of CatCoder, the benchmark datasets, and the evaluation scripts. In future work, possible research attempts include investigating the data leakage problem of the benchmarks. These investigations may help further improve the applicability of our proposed approach in practical scenarios.
Acknowledgements.
This research is supported by the Fundamental Research Funds for the Central Universities (No. 226-2025-00171).